### v1

# A dimensional mass transference principle from balls to open sets and applications to dynamical Diophantine approximation

Abstract.
The mass transference principle of Beresnevich and Velani is a powerful mechanism for determining the Hausdorff dimension/measure of $\limsup$ sets that arise naturally in Diophantine approximation. However, in the setting of dynamical Diophantine approximation, this principle often fails to apply effectively, as the radii of the balls defining the dynamical $\limsup$ sets generally depend on the orbit of the point $x$ itself.
In this paper, we develop a dimensional mass transference principle that enables us to recover and extend classical results on shrinking target problems, particularly for the $\beta$-transformation and the Gauss map. Moreover, our result shows that the corresponding $\limsup$ sets have large intersection properties. A potentially interesting feature of our method is that, in many cases, shrinking target problems are closely related to finding an appropriate Gibbs measure, which may reveal new aspects of the link between thermodynamic formalism and dynamical Diophantine approximation.
Key words and phrases:
mass transference principle, dynamical Diophantine approximation, Hausdorff dimension, Gibbs measure
2020 Mathematics Subject Classification:
11J83, 37D35
1. Introduction
The central question in Diophantine approximation is: how well can a given real number $x\in[0,1)$ be approximated by rational numbers. Dating back to Dirichlet, a consequnce of his famous theorem is that for any $x\in[0,1)$,
(1.1) $\bigg{|}x-\frac{p}{q}\bigg{|}<\frac{1}{q^{2}}\quad\text{for i.m. $\frac{p}{q}\in\mathbb{Q}$},$
where i.m. stands for infinitely many. The estimate above provides an approximation rate valid for all $x$ and lays the foundation for the metric theory of Diophantine approximation. This theory seeks to understand the sets of $x$ for which inequalities analogous to (1.1) hold, but with the right-hand-side replaced by functions of $q$ that decay more rapidly. For any $\tau\geq 2$, define
$W(\tau):=\bigg{\{}x\in[0,1):\bigg{|}x-\frac{p}{q}\bigg{|}<\frac{1}{q^{\tau}}\text{ for i.m. $\frac{p}{q}\in\mathbb{Q}$}\bigg{\}}.$
A classical result, proved independently by Besicovitch [8] and Jarník [25], shows that for any $\tau\geq 2$,
(1.2) $\operatorname{\dim_{H}}W(\tau)=2/\tau,$
where $\operatorname{\dim_{H}}$ denotes the Hausdorff dimension. Remarkably, a profound connection between the statements described in (1.1) and (1.2) was uncovered by Beresnevich and Velani [7] through their celebrated mass transference principle, a powerful tool for deriving lower bounds on the Hausdorff dimension of a broad class of $\limsup$ sets. More specifically, Dirichlet theorem alone suffices to deduce the Besicovitch–Jarník theorem via their principle. We begin with some notation before stating their principle.
Throughout, the symbols $\ll$ and $\gg$ will be used to indicate an inequality with an unspecified positive multiplicative constant. If $a\ll b$ and $a\gg b$, we write $a\asymp b$ and say that the quantities $a$ and $b$ are comparable. Let $X$ be a compact metric space equipped with a non-atomic probability measure $\mu$. Suppose there exists a constant $\delta>0$ such that
$\mu(B(x,r))\asymp r^{\delta},$
where the implied constant does not depend on $x$ and $r$. Such a measure is said to be $\delta$-Ahlfors regular.
The following statement is a simplified and slightly reformulated version of the result in [7], adapted for our purposes.
Theorem 1.1 (Mass transference principle [7, Theorem 3]).
Let $X$ be a compact metric space equipped with a $\delta$-Ahlfors regular measure $\mu$. Let $\{B(x_{n},r_{n})\}$ be a sequence of balls in $X$ with $r_{n}\to 0$ as $n\to\infty$. Suppose that
$\mu\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n})\Big{)}=1.$
Then, for any $\tau>1$,
$\operatorname{\dim_{H}}\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n}^{\tau})\Big{)}\geq\frac{\delta}{\tau}.$
The mass transference principle in this form concerns $\limsup$ sets defined by balls, which is sufficient for many classical applications. However, many naturally occurring $\limsup$ sets in Diophantine approximation are defined in terms of rectangles, neighborhoods of resonant sets, or more general open sets. To address such cases, various extensions of the mass transference principle have been developed, allowing for $\limsup$ sets defined by a wider range of shapes. We refer the reader to [3, 1, 29, 43, 39, 30, 20, 12, 13, 15, 35] for further details.
Classical Diophantine approximation concerns the distribution of rational approximations to real numbers. In recent years, this classical viewpoint has been naturally extended to the setting of dynamical Diophantine approximation, which studies approximation properties along orbits of dynamical systems. Among various problems in this field, our primary focus is on the shrinking target problem, first introduced by Hill and Velani [21], along with its generalizations, which concern whether the orbit of a given point hits a sequence of shrinking targets infinitely often.
Let $(X,d,T)$ be a dynamical system. The shrinking target problem studies the size, expressed in terms of dimension and measure, of the shrinking target set
$\{x\in X:d(T^{n}x,x_{0})<\psi(n,x)\text{ for i.m. $n$}\},$
where $x_{0}\in X$ and $\psi:\mathbb{N}\times X\to\mathbb{R}_{\geq 0}$ is a positive function. Numerous results on the measure and dimension of shrinking target sets have been established in various dynamical systems; see, for example, [2, 4, 5, 11, 19, 23, 28, 31, 36, 37, 42]. To illustrate, consider the doubling map $T_{2}(x)=2x\ (\textrm{mod}\ 1)$ on $[0,1)$. In this setting, we are interested in the shrinking target set
$W(T_{2},f,x_{0}):=\left\{x\in[0,1):|T_{2}^{n}x-x_{0}|<e^{-S_{n}f(x)}\ \text{for i.m. }n\right\},$
where $f:[0,1)\to\mathbb{R}$ is a positive function and
$S_{n}f(x)=\sum_{k=0}^{n-1}f(T_{2}^{k}x)$
is the Birkhoff sum of $f$ along the orbit of $x$. It is well-known that
$\operatorname{\dim_{H}}W(T_{2},f,x_{0})=s,$
where $s$ satisfies $P(-s(f+\log 2),T_{2})=0$. Here, $P(\cdot,T_{2})$ denotes the pressure function, see (4.1) for the definition.
A natural question is whether the mass transference principle stated in Theorem 1.1 can be applied to obtain the Hausdorff dimension of $W(T_{2},f,x_{0})$. While the principle is applicable in this setting, it may fail to yield the desired lower bound. This limitation arises because, unlike in classical Diophantine approximation, the targets here are dynamically defined and their radii depend on the orbit of $x$ itself. To apply the principle effectively, one would need to enlarge these balls by a power significantly larger than expected to obtain a $\limsup$ set with full measure. This is precisely why the classical mass transference principle does not directly provide the desired dimension estimates.
To address this shortfall, Wang and Zhang [41] developed an alternative mass transference principle from a dynamical perspective. Utilizing their principle, they successfully recovered the Hausdorff dimension of $W(T_{2},f,x_{0})$. Although the dimension result for $W(T_{2},f,x_{0})$ has been known, their work is significant in providing a new framework that connects shrinking target problems with mass transference principle. However, their principle does not extend to more general transformations such as the $\beta$-transformation or the Gauss map, nor can it be applied to settings where targets are defined by arbitrary open sets rather than balls. The main goal of this paper is to address precisely this issue. Our purpose is to develop a framework capable of handling shrinking target problems — along with various generalizations — for the $\beta$-transformation and the Gauss map, and to extend the theory beyond the classical setting of balls to more general open sets.
Part of the inspiration for our approach originates from the work of Barral and Seuret [6] and Daviaud [12], who established that for a quasi-Bernoulli probability measure $\nu$ (see, e.g., [12, Definition 2.3]),
(1.3) $\begin{split}&\nu\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n})\Big{)}=1\\ \Longrightarrow\quad&\operatorname{\dim_{H}}\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n}^{\tau})\Big{)}\geq\frac{\operatorname{\dim_{H}}\nu}{\tau}\quad\text{ for $\tau>1$}.\end{split}$
Here, the Hausdorff dimension of a measure $\nu$ is defined via its lower local dimension at $x$,
$\underline{D}(\nu,x):=\liminf_{r\to 0}\frac{\log\nu(B(x,r))}{\log r},$
and the lower and upper Hausdorff dimensions of $\nu$ are given by
[…]
If $\operatorname{\underline{\dim}_{H}}\nu=\operatorname{\overline{\dim}_{H}}\nu$, their common value is denoted by $\operatorname{\dim_{H}}\nu$. However, the quasi-Bernoulli property holds for the Gibbs measures associated with the doubling map, but generally fails for the $\beta$-transformation and the Gauss map. This limitation motivates the development of new concepts capable of handling these cases. To this end, we introduce the notion of quasi-self-conformality of a measure.
Definition 1.2 (Quasi-self-conformality).
Let $\nu$ be a Borel probability measure supported on a metric space $X$, and let $\mathcal{F}=\{F_{n}\}$ be a collection of closed subsets of $X$. We say that $\nu$ is quasi-self-conformal with respect to $\mathcal{F}$ if there exists a constant $C\geq 1$ such that for every $F_{n}\in\mathcal{F}$, there exists a bijection $f_{n}:F_{n}\to X$ satisfying:
  1. (1)
    $C^{-1}\dfrac{|x-y|}{|F_{n}|}\leq|f_{n}(x)-f_{n}(y)|\leq C\dfrac{|x-y|}{|F_{n}|}$ for all $x,y\in F_{n}$, where $|A|$ denotes the diameter of a set $A$;
  2. (2)
    The normalized pushforward measure $\nu^{(n)}:=\dfrac{\nu\circ f_{n}^{-1}}{\nu(F_{n})}$ satisfies
    $C^{-1}\nu(A)\leq\nu^{(n)}(A)\leq C\nu(A)\quad\text{for any Borel set $A$}.$
The notion of quasi-self-conformality arises as an appropriate generalization of the classical concept of self-conformality for sets, designed to capture approximately self-conformal structures exhibited by measures.
Remark 1.
Definition 1.2 (1) implies that:
  1. (a)
    For any $x\in F_{n}$ and $0<r<|F_{n}|$,
    $B(f_{n}(x),C^{-1}r/|F_{n}|)\subset f_{n}(B(x,r)\cap F_{n})\subset B(f_{n}(x),Cr/|F_{n}|).$
  2. (b)
    If $X$ supports a $\delta$-Ahlfors regular measure $\mu$, then there exists an absolute constant $c\geq 1$ such that for any $F_{n}$,
    $F_{n}\subset B(x_{n},c|F_{n}|)\mathrm{\quad and\quad}c^{-1}|F_{n}|^{\delta}\leq\mu(F_{n})\leq c|F_{n}|^{\delta},$
    where $x_{n}\in F_{n}$.
To formulate our main result, we recall the notion of Hausdorff content. In this paper, we focus on the case where the ambient space $X$ is a compact subset of $\mathbb{R}^{d}$. For any $s\geq 0$ and a set $A$, the $s$-dimensional Hausdorff content of $A$ is defined by
$\mathcal{H}_{\infty}^{s}(A)=\inf\bigg{\{}\sum_{i}|B_{i}|^{s}:A\subset\bigcup_{i\geq 1}B_{i},\text{ where $B_{i}$ are balls}\bigg{\}}.$
Our method further enables us to establish the so-called large intersection property, introduced and systematically studied by Falconer [16].
Definition 1.3 ([16]).
Let $0<s\leq\operatorname{\dim_{H}}X$. We define $\mathscr{G}^{s}(X)$ to be the class of $G_{\delta}$-subsets $A$ of $X$ such that there exists a constant $c>0$ such that for any $0<t<s$ and any ball $B$,
(1.4) $\mathcal{H}_{\infty}^{t}(A\cap B)>c\mathcal{H}_{\infty}^{t}(B).$
If $X$ supports a $\delta$-Ahlfors regular measure, then the class $\mathscr{G}^{s}(X)$ is closed under countable intersections, and moreover,
$\operatorname{\dim_{H}}A\geq s\quad\text{for all $A\in\mathscr{G}^{s}(X)$}.$
Theorem 1.4.
Let $X\subset\mathbb{R}^{d}$ be a compact subset equipped with a $\delta$-Ahlfors regular measure $\mu$. Let $\nu$ be a quasi-self-conformal measure with respect to a collection of closed sets $\mathcal{F}=\{F_{n}\}$ in $X$, such that
(1.5) $\mu\Big{(}\limsup_{n\to\infty}F_{n}\Big{)}=1.$
Suppose that there exist a sequence of balls $\{B(x_{n},r_{n})\}$ and a sequence of open sets $\{E_{n}\}$ satisfying the following conditions:
  1. (1)
    $r_{n}\to 0$ as $n\to\infty$;
  2. (2)
    $\mu\big{(}\limsup B(x_{n},r_{n})\big{)}=1$;
  3. (3)
    $E_{n}\subset B_{n}$. Moreover, there exists a constant $s\geq 0$ such that
    $\mathcal{H}_{\infty}^{s}(E_{n})\gg r_{n}^{\operatorname{\overline{\dim}_{H}}\mu},$
    where the implied constant is independent of $n$.
Then,
$\limsup_{n\to\infty}E_{n}\in\mathscr{G}^{s}(X).$
Remark 2.
The assumption that $X\subset\mathbb{R}^{d}$ is essential, as our arguments rely on the Besicovitch covering theorem, which generally does not hold in metric spaces. This arises because the measure $\nu$ is generally not doubling, which prevents many standard covering lemmas from applying effectively — except for the Besicovitch covering theorem. The condition in (1.5) serves to characterize the extent to which the measure $\nu$ exhibits quasi-self-conformality. In many case, a measure $\nu$ satisfy this condition is a Gibbs measure, and thus generally singular to the ambient measure $\mu$.
There are two different but closely related ways to understand the connection between our result and the mass transference principle of Beresnevich and Velani. To set the stage, suppose there exist two sequences of balls $\{B(x_{n},r_{n})\}$ and $\{B(y_{n},t_{n})\}$ such that
$\mu\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n})\Big{)}=1\mathrm{\quad and\quad}\nu\Big{(}\limsup_{n\to\infty}B(y_{n},t_{n})\Big{)}=1,$
where $\mu$ and $\nu$ are measures as in Theorem 1.4. The first perspective may be regarded as a counterpart to (1.3). Fix $\tau>1$, and apply Theorem 1.1 and Theorem 1.4 with $s=\operatorname{\overline{\dim}_{H}}\nu/\tau$, respectively. Note that $\delta=\operatorname{\dim_{H}}\mu$. Then, we obtain
$\operatorname{\dim_{H}}\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n}^{\tau})\Big{)}\geq\operatorname{\dim_{H}}\mu/\tau\mathrm{\quad and\quad}\operatorname{\dim_{H}}\Big{(}\limsup_{n\to\infty}B(y_{n},t_{n}^{\tau})\Big{)}\geq\operatorname{\overline{\dim}_{H}}\nu/\tau,$
and these lower bounds coincide when $\mu=\nu$. In other words, if enlarging the radii of the balls by a power of $1/\tau$ yields a $\limsup$ set of full measure (with respect to $\mu$ or $\nu$), then the $\limsup$ set of the original balls has Hausdorff dimension at least
$\frac{\text{Hausdorff dimension of the measure}}{\tau}.$
The second perspective is expressed as follows:
$\begin{split}&\mathcal{H}_{\infty}^{\delta/\tau}(B(x_{n},r_{n}^{\tau}))\asymp r_{n}^{\delta}\asymp\mu(B(x_{n},r_{n}))\quad\text{and}\quad\mu\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n})\Big{)}=1\\ \Longrightarrow\quad&\operatorname{\dim_{H}}\Big{(}\limsup_{n\to\infty}B(x_{n},r_{n}^{\tau})\Big{)}\geq\delta/\tau,\end{split}$
while
$\begin{split}&\mathcal{H}_{\infty}^{\operatorname{\overline{\dim}_{H}}\nu/\tau}(B(y_{n},t_{n}^{\tau}))\asymp r_{n}^{\operatorname{\overline{\dim}_{H}}\nu}\sim\nu(B(y_{n},t_{n}))\quad\text{and}\quad\nu\Big{(}\limsup_{n\to\infty}B(y_{n},t_{n})\Big{)}=1\\ \Longrightarrow\quad&\operatorname{\dim_{H}}\Big{(}\limsup_{n\to\infty}B(y_{n},t_{n}^{\tau})\Big{)}\geq\operatorname{\overline{\dim}_{H}}\nu/\tau.\end{split}$
Here we use $\sim$’ instead of $\asymp$’ because two quantities in general no comparable. That is, if a $\limsup$ set has full measure (with respect to $\mu$ or $\nu$) and we shrink each ball to a smaller one whose $s$-dimensional Hausdorff content is comparable to the measure of the original, then the resulting $\limsup$ set has Hausdorff dimension at least $s$. Although the two perspectives are equivalent, the latter is more flexible and naturally leads us to consider extending from balls to general open sets.
We now apply Theorem 1.4 to recover and extend classical results for the $\beta$-transformation and the Gauss map. For a Lipschitz function $h$, its Lipschitz constant is defined as the smallest $L>0$ such that for any $x,y\in X$,
$|h(x)-h(y)|\leq L|x-y|.$
Let $\{h_{n}\}$ be a sequence of Lipschitz functions with uniformly bounded Lipschitz constants and let $f$ be a positive function defined on $X$. The modified shrinking target sets are defined as
$W(T,f,\{h_{n}\})=\{x\in X:|T^{n}x-h_{n}(x)|<e^{-S_{n}f(x)}\text{ for i.m. $n$}\}.$
Theorem 1.5.
Suppose that $T$ is either the $\beta$-transformation or the Gauss map. Then,
$W(T,f,\{h_{n}\})\in\mathscr{G}^{s}([0,1]),$
where $s$ is the unique solution to $P(-s(f+\log|T^{\prime}|),T)=0$.
Remark 3.
The lower bound for the Hausdorff dimension of $W(T,f,\{h_{n}\})$ implied in Theorem 1.5 was previously established in [10, 37, 32, 42]. However, those results rely on the construction of large Cantor-type sets and do not imply the large intersection property. Interestingly, Theorem 1.4 offers a different perspective: the problem is reduced to seeking a suitable Gibbs measure and estimating its Hausdorff dimension. This perspective may provide new insights into the interplay between thermodynamic formalism and dynamical Diophantine approximation.
Remark 4.
After completing the proofs of our main results, the author became aware that Daviaud [14] had employed some similar ideas to study the shrinking target problem for self-conformal sets with overlaps. However, his results neither imply the large intersection property nor can they be directly applied to the $\beta$-transformation or the Gauss map.
Theorem 1.5 is a direct application of Theorem 1.4, where the sets $E_{n}$ are taken to be balls. To further demonstrate the versatility of our main result, we present a concise proof of the following theorem, which was also previously established in [22]. Let $m\geq 1$ be an integer and $B>1$. Define
$F_{m}(B):=\{x\in[0,1):a_{n+1}(x)\cdots a_{n+m}(x)\geq B^{n}\text{ for i.m. $n$}\},$
where $a_{n}(x)$ denotes the $n$th partial quotient of $x$ (see Section 5 for the definition).
Theorem 1.6.
Let $m\geq 1$ be an integer and $B>1$. Then,
$F_{m}(B)\in\mathscr{G}^{u}([0,1]),$
for some $u\in(1/2,1)$ satisfying
$P(-u\log|G^{\prime}|-g_{m}(u)\log B,G)=0,$
where the function $g_{m}(u)$ is given by
$g_{m}(u)=\frac{u^{m}(2u-1)}{u^{m}-(1-u)^{m}}.$
The structure of the paper is as follows. In Section 2, we collect several foundational results and technical tools that will be used throughout the paper. Section 3 is devoted to the proof of our main result, i.e. Theorem 1.4. In Section 4, we review the definition and key properties of the $\beta$-transformation, and apply our main result to obtain the large intersection property of $W(T_{\beta},f,\{h_{n}\})$ in this setting. Section 5 serves a similar purpose for the Gauss map: we recall its basic properties and then apply our theorem to derive the large intersection properties of $W(G,f,\{h_{n}\})$ and $F_{m}(B)$.
2. Preliminary
This section recalls key tools from geometric measure theory and covering arguments that underpin the main results of this paper. We start by presenting two fundamental results that relate the Hausdorff content of a set to probability measures exhibiting appropriate local dimension estimates. Here and hereafter, we will assume that $X\subset\mathbb{R}^{d}$ is compact equipped with a $\delta$-Ahlfors regular measure $\mu$.
Proposition 2.1 (Mass distribution principle [9, Lemma 1.2.8]).
Let $A$ be a Borel subset of $\mathbb{R}^{d}$. If $A$ supports a Borel probability measure $\lambda$ that satisfies
$\lambda(B(x,r))\leq cr^{s},$
for some constant $0<c<\infty$, and for every ball $x\in\mathbb{R}^{d}$ and $r>0$, then
$\mathcal{H}_{\infty}^{s}(A)\geq 1/c.$
Lemma 2.2 (Frostman’s lemma [33, Theorem 8.8]).
Let $A$ be a Borel subset of $\mathbb{R}^{d}$. If $\mathcal{H}_{\infty}^{s}(A)>0$, then there exists a probability measure $\lambda$ supported on $A$ such that for any $x\in\mathbb{R}^{d}$ and $r>0$,
$\lambda(B(x,r))\ll\frac{r^{s}}{\mathcal{H}_{\infty}^{s}(A)},$
where the unspecified constant depends only on $d$.
Theorem 2.3 offers a relatively simple criterion for verifying that a $\limsup$ set has the large intersection property.
Theorem 2.3 ([20, Corollary 2.6]).
Let $0<s\leq\operatorname{\dim_{H}}X$. Let $\{E_{n}\}$ be a sequence of open sets in $X$. If for any $0<t<s$, there exists a constant $c=c(t)>0$ such that
$\limsup_{n\to\infty}\mathcal{H}_{\infty}^{t}(E_{n}\cap B)>c\mu(B)$
holds for any ball $B\subset X$, then
$\limsup_{n\to\infty}E_{n}\in\mathscr{G}^{s}(X).$
The following covering result, due to Besicovitch, will be used to efficiently select disjoint subfamilies of balls covering a given set.
Theorem 2.4 (Besicovitch covering Theorem [33, Theorem 2.7]).
There is a positive integer $Q_{d}$ depending only on the dimension $d$ with the following property. Let $A\subset\mathbb{R}^{d}$ be a bounded set, and let $\mathcal{B}$ be a family of balls such that each point of $A$ is the centre of some ball of $\mathcal{B}$. There are families $\mathcal{B}_{1},\dots,\mathcal{B}_{Q_{d}}\subset\mathcal{B}$ covering $A$ such that each $\mathcal{B}_{k}$ is disjoint, that is,
$A\subset\bigcup_{1\leq k\leq Q_{d}}\bigcup_{B\in\mathcal{B}_{k}}B$
and
[…]
The next lemma allows us to extract well-separated subcollections from a sequence of shrinking balls while retaining a definite portion of total measure.
Lemma 2.5 ([7, Lemma 5]).
Let $\{B(x_{n},r_{n})\}$ be a sequence of balls in $X\subset\mathbb{R}^{d}$ with $r_{n}\to 0$ as $n\to\infty$. Then, for any ball $B$ in $X$, there exists a finite collection
$K_{B}\subset\{B(x_{n},r_{n})\}$
satisfying the following properties:
  1. (1)
    $B(x_{n},r_{n})\subset B$ for all $B(x_{n},r_{n})\in K_{B}$;
  2. (2)
    If $B(x_{n},r_{n}),B(x_{m},r_{m})\in K_{B}$ are distinct, then $B(x_{n},3r_{n})\cap B(x_{m},3r_{m})=\emptyset$;
  3. (3)
    there exists a constant $c>0$ independent of $B$ such that
    $\mu\bigg{(}\bigcup_{B(x_{n},r_{n})\in K_{B}}B(x_{n},r_{n})\bigg{)}\geq c\mu(B).$
Here, we highlight the difference between Besicovitch covering theorem and Lemma 2.5. Besicovitch covering theorem applies to arbitrary measures but is restricted to Euclidean spaces, whereas Lemma 2.5 can be extended to general metric spaces but requires the measure to be doubling. In the sequel, when it is necessary to extract a disjoint subcollection from a sequence of shrinking balls, we will use Lemma 2.5 for the ambient measure $\mu$, and the Besicovitch covering theorem for the reference measure $\nu$.
Note that as the collection $\{F_{n}\}$ of closed sets in Theorem 1.4 are not necessarily balls, the following variant of Lemma 2.5 is needed.
Corollary 2.6.
Let $\mathcal{F}=\{F_{n}\}$ be as given in Theorem 1.4. That is, each $F_{n}\in\mathcal{F}$ satisfies Definition 1.2 (1) and
$\mu\Big{(}\limsup_{n\to\infty}F_{n}\Big{)}=1.$
Then, for any ball $B$ in $X$, there exists a finite collection
$\mathcal{F}_{B}\subset\mathcal{F}$
satisfying the following properties:
  1. (1)
    $F_{n}\subset B$ for all $F_{n}\in\mathcal{F}_{B}$;
  2. (2)
    If $F_{n}$ and $F_{m}$ are distinct then $\mathrm{dist}(F_{n},F_{m})\geq\max\{|F_{n}|,|F_{n^{\prime}}|\}$, where the distance between sets is defined by
    $\mathrm{dist}(F_{n},F_{m}):=\inf\{d(x,y):x\in F_{n},\ y\in F_{m}\};$
  3. (3)
    there exists a constant $c^{\prime}>0$ independent of $B$ such that
    $\mu\bigg{(}\bigcup_{F_{n}\in\mathcal{F}_{B}}F_{n}\bigg{)}\geq c^{\prime}\mu(B).$
Proof.
Recall from Remark 1 (b) that there exists an absolute constant $c\geq 1$ such that for any $F_{n}$,
(2.1) $F_{n}\subset B(x_{n},c|F_{n}|)\mathrm{\quad and\quad}c^{-1}|F_{n}|^{\delta}\leq\mu(F_{n})\leq c|F_{n}|^{\delta},$
where $x_{n}\in F_{n}$. Clearly,
$\mu\Big{(}\limsup_{n\to\infty}B(x_{n},c|F_{n}|)\Big{)}=1$
Applying Lemma 2.5 to the sequence $\{B(x_{n},c|F_{n}|)\}$, we obtain a finite collection $K_{B}$ satisfying items (1)–(3) in Lemma 2.5. Let
$\mathcal{F}_{B}=\{F_{n}:B(x_{n},c|F_{n}|)\in K_{B}\}.$
By (2.1), items (1) and (3) in corollary follows immediately from those in Lemma 2.5.
For item (2), suppose that $F_{n},F_{m}\in\mathcal{F}_{B}$ are distinct. Then, by definition, the same is true for $B(x_{n},c|F_{n}|),B(x_{m},c|F_{m}|)\in K_{B}$. It follows that
$B(x_{n},3c|F_{n}|)\cap B(x_{m},3c|F_{m}|)=\emptyset.$
Therefore,
$\mathrm{dist}\big{(}B(x_{n},c|F_{n}|),B(x_{m},c|F_{m}|)\big{)}\geq\max\{c|F_{n}|,c|F_{m}|\}\geq\max\{|F_{n}|,|F_{m}|\}.$
Consequently,
$\mathrm{dist}(F_{n},F_{m})\geq\mathrm{dist}\big{(}B(x_{n},c|F_{n}|),B(x_{m},c|F_{m}|)\big{)}\geq\max\{|F_{n}|,|F_{m}|\}.\qed$
3. Proof of Theorem 1.4
Let $\{E_{n}\}$ be as given in Theorem 1.4. In this section, our goal is to establish that the corresponding $\limsup$ set has the large intersection property. Specifically, we will show that for any $0<t<s$,
$\limsup_{\ell\to\infty}\mathcal{H}_{\infty}^{t}\bigg{(}\bigcup_{k=\ell}^{\infty}E_{k}\cap B\bigg{)}\gg\mu(B)\quad\text{holds for all ball $B\subset X$},$
where the implied constant is independent of the ball $B$. Once this is established, Theorem 2.3 yields
$\limsup_{\ell\to\infty}\bigg{(}\bigcup_{k=\ell}^{\infty}E_{k}\bigg{)}=\limsup_{n\to\infty}E_{n}\in\mathscr{G}^{s}(X),$
thereby completing the proof of the large intersection property for the set $\limsup E_{n}$.
To proceed, fix $0<t<s$, $\ell\geq 1$, and a ball $B\subset X$. The remainder of this section is devoted to establishing the lower bound
(3.1) $\mathcal{H}_{\infty}^{t}\bigg{(}\bigcup_{k=\ell}^{\infty}E_{k}\cap B\bigg{)}\gg\mu(B).$
3.1. Construction of a subset of $\bigcup_{k=\ell}^{\infty}E_{k}\cap B$
Our approach to constructing the desired subset is motivated by [12], but the overall strategy we adopt to establish the lower bound of the Hausdorff dimension of $\limsup E_{n}$ is different.
Let $\nu$ be the reference measure stated in Theorem 1.4. Let $\varepsilon=s-t>0$. To make effective use of the local behavior of the measure $\nu$, we consider the set of points where the lower local dimension exceeds a certain threshold. By the definition of the lower Hausdorff dimension $\operatorname{\overline{\dim}_{H}}\nu$, the set
$E_{\nu}^{\varepsilon}:=\{x\in X:\underline{D}(\nu,x)>\operatorname{\overline{\dim}_{H}}\nu-\varepsilon\}$
has positive $\nu$-measure. Let us denote $\gamma_{\varepsilon}:=\nu(E_{\nu}^{\varepsilon})$ for convenience. To obtain a uniform estimate on the measure of small balls, we consider the sets
$E_{\nu}^{n,\varepsilon}:=\{x\in X:\text{$\forall$ $0<r<1/n$, $\nu(B(x,r))\leq r^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}$}\}.$
By definition, we have
$E_{\nu}^{\varepsilon}=\bigcup_{n=1}^{\infty}E_{\nu}^{n,\varepsilon},$
and clearly the sequence $\{E_{\nu}^{n,\varepsilon}\}$ is increasing in $n$. Therefore, by the continuity of measure from below, there exists an integer $N=N(\varepsilon)$ such that
(3.2) $\nu(E_{\nu}^{N,\varepsilon})\geq\gamma_{\varepsilon}/2.$
For the given ball $B$, let $\mathcal{F}_{B}\subset\mathcal{F}$ be the finite subcollection of balls obtained from Corollary 2.6. Recall that for any $F_{i}\subset\mathcal{F}$,
(3.3) $\nu/C\leq\nu^{(i)}=\frac{\nu\circ f_{i}^{-1}}{\nu(F_{i})}\leq C\nu,$
where $f_{i}:F_{i}\to X$ is a bijection. Let $F_{i}\in\mathcal{F}_{B}$. For any $x\in F_{i}$ and $0<\rho<|F_{i}|$, it follows from Definition 1.2 that $f_{i}$ is injective on $B(x,\rho)\cap F_{i}$. Moreover, we estimate the measure of a small ball intersected with $F_{i}$ as follows:
$\nu(B(x,\rho)\cap F_{i})$ $=\nu\big{(}f_{i}^{-1}(f_{i}(B(x,\rho)\cap F_{i}))\big{)}=\nu(F_{i})\nu^{(i)}\big{(}f_{i}(B(x,\rho)\cap F_{i})\big{)}$
(3.4) $\leq C\nu(F_{i})\nu\big{(}B(f_{i}(x),C\rho/|F_{i}|)\big{)},$
where the last inequality follows from (3.3) and item (1) of Remark 1.
Let $x\in f_{i}^{-1}(E_{\nu}^{N,\varepsilon})$. Then, $x\in F_{i}$ and $f_{i}(x)\in E_{\nu}^{N,\varepsilon}$. By the definition of $E_{\nu}^{N,\varepsilon}$, for any $0<r<1/N$,
(3.5) $\nu(B(f_{i}(x),r))\leq r^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}.$
Then, for any $0<\rho<|F_{i}|/(CN)$ (or equivalently $0<C\rho/|F_{i}|<1/N$), applying (3.4) and the above inequality yields
$\begin{split}\nu(B(x,\rho)\cap F_{i})&\leq C\nu(F_{i})\nu\bigg{(}B\bigg{(}f_{i}(x),\frac{C\rho}{|F_{i}|}\bigg{)}\bigg{)}\leq C\nu(F_{i})\cdot\bigg{(}\frac{C\rho}{|F_{i}|}\bigg{)}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}\\ &\leq C^{d+1}\nu(F_{i})\cdot\bigg{(}\frac{\rho}{|F_{i}|}\bigg{)}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}.\end{split}$
Equivalently, for any such $\rho$,
$\frac{\nu(B(x,\rho)\cap F_{i})}{\nu(F_{i})}\leq C^{d+1}\bigg{(}\frac{\rho}{|F_{i}|}\bigg{)}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}.$
Therefore, we conclude that
$\begin{split}&f_{i}^{-1}(E_{\nu}^{N,\varepsilon})\\ =&f_{i}^{-1}\big{(}\{x\in X:\text{$\forall$ $0<r<1/N$, $\nu(B(x,r))\leq r^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}$}\}\big{)}\\ \subset&\bigg{\{}x\in F_{i}:\text{$\forall$ $0<\rho<\frac{|F_{i}|}{CN}$, $\frac{\nu(B(x,\rho)\cap F_{i})}{\nu(F_{i})}\leq C^{d+1}\bigg{(}\frac{\rho}{|F_{i}|}\bigg{)}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}$}\bigg{\}}.\end{split}$
Define
$\begin{split}E_{\nu,F_{i}}^{N,\varepsilon}:=&\limsup_{n\to\infty}B(x_{n},r_{n})\cap\\ &\bigg{\{}x\in F_{i}:\text{$\forall$ $0<\rho<\frac{|F_{i}|}{CN}$, $\frac{\nu(B(x,\rho)\cap F_{i})}{\nu(F_{i})}\leq C^{d+1}\bigg{(}\frac{\rho}{|F_{i}|}\bigg{)}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}$}\bigg{\}}.\end{split}$
Since $\nu(\limsup B(x_{n},r_{n}))=1$, we have $f_{i}^{-1}(E_{\nu}^{N,\varepsilon})\subset E_{\nu,F_{i}}^{N,\varepsilon}$ except for a set of zero $\nu$-measure. Therefore,
$\nu(E_{\nu,F_{i}}^{N,\varepsilon})\geq\nu\big{(}f_{i}^{-1}(E_{\nu}^{N,\varepsilon})\big{)}=\nu(F_{i})\nu^{(i)}(E_{\nu}^{N,\varepsilon})\geq\frac{\nu(F_{i})\nu(E_{\nu}^{N,\varepsilon})}{C}\geq\frac{\gamma_{\varepsilon}\nu(F_{i})}{2C},$
where the last inequality follows from (3.2).
For any $z\in E_{\nu,F_{i}}^{N,\varepsilon}$, there exists infinitely many $n$ such that $z\in B(x_{n},r_{n})$. Choose an integer $n_{z}\geq\ell$ large enough so that
(3.6) $z\in B(x_{n_{z}},r_{n_{z}})\subset B\mathrm{\quad and\quad}16r_{n_{z}}\leq|F_{i}|/(CN).$
The above inclusion $B(x_{n_{z}},r_{n_{z}})\subset B$ is possible since $E_{\nu,F_{i}}^{N,\varepsilon}\subset F_{i}\subset B$ and $B$ is open. Set $L_{n_{z}}:=B(z,5r_{n_{z}})$. Then, we have
(3.7) $E_{n_{z}}\subset B(x_{n_{z}},r_{n_{z}})\subset L_{n_{z}}.$
Thus, the collection of balls $\{L_{n_{z}}:z\in E_{\nu,F_{i}}^{N,\varepsilon}\}$ forms a covering of $E_{\nu,F_{i}}^{N,\varepsilon}$. By Besicovitch covering theorem, one can extract from this cover a finite number (at most $Q_{d}$) of disjoint subcollections $\mathcal{B}_{k}(F_{i})$ for $1\leq k\leq Q_{d}$, such that:
  1. (1)
    Each collection $\mathcal{B}_{k}(F_{i})$ consists of pairwise disjoint balls: for any distinct $L_{n_{z}},L_{n_{w}}\in\mathcal{B}_{k}(F_{i})$, it holds that $L_{n_{z}}\cap L_{n_{w}}=\emptyset$;
  2. (2)
    The union of these collections covers the entire set:
    $E_{\nu,F_{i}}^{N,\varepsilon}\subset\bigcup_{1\leq k\leq Q_{d}}\bigcup_{L_{n_{z}}\in\mathcal{B}_{k}(F_{i})}L_{n_{z}}.$
Since $\nu(E_{\nu,F_{i}}^{N,\varepsilon})\geq\gamma_{\varepsilon}\nu(F_{i})/(2C)$, there exists some $1\leq k_{i}\leq Q_{d}$ such that the corresponding collection $\mathcal{B}_{k_{i}}(F_{i})$ satisfies
$\nu\bigg{(}\bigcup_{L_{n_{z}}\in\mathcal{B}_{k_{i}}(F_{i})}L_{n_{z}}\bigg{)}\geq\frac{\nu(E_{\nu,F_{i}}^{N,\varepsilon})}{Q_{d}}\geq\frac{\gamma_{\varepsilon}\nu(F_{i})}{2CQ_{d}}.$
By the disjointness of the balls in $\mathcal{B}_{k_{i}}(F_{i})$, we may further extract a finite subcollection $\mathcal{B}(F_{i})\subset\mathcal{B}_{k_{i}}(F_{i})$ such that
(3.8) $\nu\bigg{(}\bigcup_{L_{n_{z}}\in\mathcal{B}(F_{i})}L_{n_{z}}\bigg{)}\geq\nu\bigg{(}\bigcup_{L_{n_{z}}\in\mathcal{B}_{k_{i}}(F_{i})}L_{n_{z}}\bigg{)}\bigg{/}2\geq\frac{\gamma_{\varepsilon}\nu(F_{i})}{4CQ_{d}}.$
Note that from (3.7), each $E_{n_{z}}\subset L_{n_{z}}$, so the union
(3.9) $A:=\bigcup_{F_{i}\subset\mathcal{F}_{B}}\bigcup_{L_{n_{z}}\in\mathcal{B}(F_{i})}E_{n_{z}}\subset\bigcup_{k=\ell}^{\infty}E_{k}\cap B$
is a subset of the relevant tail of the $\limsup$ set intersected with the ball $B$.
In the next subsection, we will construct a probability measure supported on the set $A$, and show that $\mathcal{H}_{\infty}^{t}(A)\gg\mu(B)$. This will immediately yield the desired lower bound in (3.1), completing the proof of the large intersection property. Before moving to this task, we summarize several geometric and measure-theoretic properties of $A$ established so far. These will be instrumental in the measure construction and content estimates that follow.
Lemma 3.1.
Let $A$ be the set defined in (3.9). Then the following properties hold:
  1. (1)
    We have the lower bound
    (3.10) $\sum_{F_{i}\in\mathcal{F}_{B}}\mu(F_{i})\gg\mu(B).$
    Furthermore, for any two distinct sets $F_{i},F_{j}\in\mathcal{F}_{B}$,
    $\mathrm{dist}(F_{i},F_{j})\geq\max\{|F_{i}|,|F_{j}|\}.$
  2. (2)
    For each $F_{i}\in\mathcal{F}_{B}$, we have
    (3.11) $\nu\bigg{(}\bigcup_{L_{n_{z}}\in\mathcal{B}(F_{i})}L_{n_{z}}\bigg{)}\gg\nu(F_{i}),$
    where $L_{n_{z}}=B(z,5r_{n_{z}})$ is a ball with center $z\in E_{\nu,F_{i}}^{N,\varepsilon}$. Moreover, for any two distinct balls $L_{n_{z}},L_{n_{w}}\in\mathcal{B}(F_{i})$, we have
    (3.12) $L_{n_{z}}\cap L_{n_{w}}=\emptyset,\quad\text{and}\quad\mathrm{dist}(E_{n_{z}},E_{n_{w}})\geq\max\{r_{n_{z}},r_{n_{w}}\}.$
Proof.
(1) It follows from immediately from Corollary 2.6.
(2) Equation (3.11) is just a reformulation of (3.8). By the construction of the collection $\mathcal{B}(F_{i})$, the balls it contains are pairwise disjoint. Thus, it remains to verify the separation property:
$\mathrm{dist}(E_{n_{z}},E_{n_{w}})\geq\max\{r_{n_{z}},r_{n_{w}}\}.$
This follows from two observations (see (3.6) and (3.7)): first, the center $z$ lies in $B(x_{n_{z}},r_{n_{z}})$; second, the set $E_{n_{z}}$ is contained in $B(x_{n_{z}},r_{n_{z}})$, which in turn is contained in $L_{n_{z}}=B(z,5r_{n_{z}})$. These nested inclusions guarantee that the sets $E_{n_{z}}$ are mutually disjoint and separated by at least $\max\{r_{n_{z}},r_{n_{w}}\}$, since they are contained in disjoint balls $L_{n_{z}}$. ∎
3.2. Hausdorff content bound of $\bigcup_{k=\ell}^{\infty}E_{k}\cap B$
Recall condition (3) in Theorem 1.4,
$\mathcal{H}_{\infty}^{s}(E_{n})\gg r_{n}^{\operatorname{\overline{\dim}_{H}}\nu}.$
For any $n\geq 1$, by Frostman’s lemma, there exists a probability measure $\lambda_{n}$ supported on $E_{n}$ such that
(3.13) $\lambda_{n}(B(x,r))\ll\frac{r^{s}}{r_{n}^{\operatorname{\overline{\dim}_{H}}\nu}}.$
Since $E_{n}\subset B(x_{n},r_{n})$, it follows that
$s\leq\operatorname{\overline{\dim}_{H}}\nu\leq\delta.$
Let $A$ be defined as in (3.9). Define a probability measure $\eta$ supported on $A\subset\bigcup_{k=\ell}^{\infty}E_{k}\cap B$ by
$\eta=\sum_{F_{i}\in\mathcal{F}_{B}}\sum_{L_{n_{z}}\in\mathcal{B}(F_{i})}\frac{\mu(F_{i})}{\sum_{F_{i}\subset\mathcal{F}_{B}}\mu(F_{i})}\cdot\frac{\nu(L_{n_{z}})}{\sum_{L_{n_{z}}\in\mathcal{B}(F_{i})}\nu(L_{n_{z}})}\cdot\lambda_{n_{z}}.$
Next, we estimate the $\eta$-measure of arbitrarily balls, which will allow us to apply the mass distribution principle and conclude the desired lower bound on the Hausdorff content. Suppose that $r>0$ and
(3.14) $x\in E_{n_{w}}\quad\text{for some\quad$E_{n_{w}}\subset L_{n_{w}}=B(w,5r_{n_{w}})\in\mathcal{B}(F_{i})$}.$
The separation properties of the collections $\mathcal{F}_{B}$ and $\mathcal{B}(F_{i})$ (established in Lemma 3.1) suggest us to consider four different cases.
Case 1: $r>|B|$. Since $\eta$ is a probability measure supported on a subset of $B$, the measure of any ball with radius larger than $|B|$ is trivially bounded by 1. Using the $\delta$-Ahlfors regularity of $\mu$, we have
(3.15) $\eta(B(x,r))\leq 1<\frac{r^{\delta}}{|B|^{\delta}}\ll\frac{r^{s}}{\mu(B)}.$
Case 2: $|F_{i}|\leq r<|B|$. By the separation property of the collection $\mathcal{F}_{B}$, different sets $F_{j}$ are well spaced apart. Specifically, for any $F_{j}\in\mathcal{F}_{B}$ distinct with $F_{i}$, by Lemma 3.1 (1),
(3.16) $\mathrm{dist}(F_{i},F_{j})\geq\max\{|F_{i}|,|F_{j}|\}.$
If a distinct $F_{j}$ intersects $B(x,r)$, then its diameter must be at most $r$, implying that $F_{j}$ lies within a slightly larger ball $B(x,2r)$. It follows that
(3.17) $\eta(B(x,r))\leq\sum_{\begin{subarray}{c}F_{i}\in\mathcal{F}_{B}\\ F_{i}\subset B(x,2r)\end{subarray}}\frac{\mu(F_{i})}{\sum_{F_{i}\subset\mathcal{F}_{B}}\mu(F_{i})}\ll\frac{\mu(B(x,2r))}{\mu(B)}\ll\frac{r^{\delta}}{\mu(B)}\ll\frac{r^{s}}{\mu(B)}.$
Case 3: $r_{n_{w}}\leq r<|F_{i}|$. Here, $B(x,r)$ intersects only one $F_{i}$ because of the minimal distance (see (3.16)) between distinct sets $F_{i}$. We break it down into two subcases:
Subcase 3a: $r\geq|F_{i}|/(16CN)$. By the definition of $\eta$,
(3.18) $\eta(B(x,r))\leq\frac{\mu(F_{i})}{\sum_{F_{i}\subset\mathcal{F}_{B}}\mu(F_{i})}\ll\frac{|F_{i}|^{\delta}}{\mu(B)}\leq\frac{(16CNr)^{\delta}}{\mu(B)}\ll\frac{r^{s}}{\mu(B)},$
where we use the fact that $N=N(\varepsilon)$ is independent of $B$ (see (3.2)).
Subcase 3b: $r_{n_{w}}\leq r<|F_{i}|/(16CN)$. For any ball $L_{n_{z}}\in\mathcal{B}(F_{i})$ distinct with $L_{n_{w}}$, if $B(x,r)\cap E_{n_{z}}=\emptyset$, then
$\lambda_{n_{z}}(B(x,r))=0,$
since $\lambda_{n_{z}}$ is supported on the set $E_{n_{z}}$. Consequently, we have
$\begin{split}\eta(B(x,r))&\leq\sum_{\begin{subarray}{c}L_{n_{z}}\in\mathcal{B}(F_{i})\\ B(x,r)\cap E_{n_{z}}\neq\emptyset\end{subarray}}\frac{\mu(F_{i})}{\sum_{F_{i}\subset\mathcal{F}_{B}}\mu(F_{i})}\cdot\frac{\nu(L_{n_{z}})}{\sum_{L_{n_{z}}\in\mathcal{B}(F_{i})}\nu(L_{n_{z}})}\\ &\ll\sum_{\begin{subarray}{c}L_{n_{z}}\in\mathcal{B}(F_{i})\\ B(x,r)\cap E_{n_{z}}\neq\emptyset\end{subarray}}\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(L_{n_{z}})}{\nu(F_{i})}.\end{split}$
Note that by (3.12) the sets $E_{n_{z}}$ and $E_{n_{w}}$ are well-separated:
(3.19) $\mathrm{dist}(E_{n_{z}},E_{n_{w}})\geq\max\{r_{n_{z}},r_{n_{w}}\}.$
Therefore, if $B(x,r)\cap E_{n_{z}}\neq\emptyset$, then it must be that
$\begin{split}r>r_{n_{z}}\geq|L_{n_{z}}|/10\quad\Longrightarrow\quad E_{n_{z}}\subset L_{n_{z}}\subset B(x,11r)\subset B(w,16r),\end{split}$
where the last inclusion uses the fact that $x\in E_{n_{w}}\subset B(w,5r_{n_{w}})$ (see (3.14)). It follows that
$\begin{split}\eta(B(x,r))&\ll\sum_{\begin{subarray}{c}L_{n_{z}}\in\mathcal{B}(F_{i})\\ B(x,r)\cap E_{n_{z}}\neq\emptyset\end{subarray}}\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(L_{n_{z}})}{\nu(F_{i})}.\ll\sum_{\begin{subarray}{c}L_{n_{z}}\in\mathcal{B}(F_{i})\\ L_{n_{z}}\subset B(w,16r)\end{subarray}}\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(L_{n_{z}})}{\nu(F_{i})}\\ &\leq\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(B(w,16r)\cap F_{i})}{\nu(F_{i})}\end{split}$
Note that $w\in E_{\nu,F_{i}}^{N,\varepsilon}$. Since $r<|F_{i}|/(16CN)$ (equivalently $16r<|F_{i}|/(CN)$), by the definition of $E_{\nu,F_{i}}^{N,\varepsilon}$,
$\frac{\mu(B(w,16r)\cap F_{i})}{\nu(F_{i})}\ll\bigg{(}\frac{16r}{|F_{i}|}\bigg{)}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}\ll\frac{r^{s-\varepsilon}}{|F_{i}|^{\delta}},$
where we use $s\leq\operatorname{\overline{\dim}_{H}}\nu\leq\delta$ in the last step. Putting all these together, we conclude that
(3.20) $\eta(B(x,r))\ll\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(B(w,16r)\cap F_{i})}{\nu(F_{i})}\ll\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{r^{s-\varepsilon}}{|F_{i}|^{\delta}}\ll\frac{r^{s-\varepsilon}}{\mu(B)}.$
Case 4: $0<r<r_{n_{w}}$. In this scale, due to the separation of the sets $E_{n_{z}}$ (see (3.12)), the ball $B(x,r)$ can intersect only the single set $E_{n_{w}}$ containing $x$. By the Frostman-type property for $\lambda_{n_{w}}$ (see (3.13)),
$\eta(B(x,r))$ $\ll\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(L_{n_{w}})}{\nu(F_{i})}\cdot\lambda_{n_{w}}(B(x,r))\ll\frac{|F_{i}|^{\delta}}{\mu(B)}\cdot\bigg{(}\frac{r_{n_{w}}}{|F_{i}|}\bigg{)}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}\cdot\frac{r^{s}}{r_{n_{w}}^{\operatorname{\overline{\dim}_{H}}\nu}}$
(3.21) $\ll\frac{r_{n_{w}}^{-\varepsilon}r^{s}}{\mu(B)}\leq\frac{r^{s-\varepsilon}}{\mu(B)}.$
By Cases 1–4, we have for any ball $B(x,r)$,
$\eta(B(x,r))\ll\frac{r^{s-\varepsilon}}{\mu(B)}=\frac{r^{t}}{\mu(B)},$
where the equality follows from $\varepsilon=s-t$. Since $\eta$ is supported on $A\subset\bigcup_{k=\ell}^{\infty}E_{k}\cap B$, by the mass distribution principle,
$\mathcal{H}_{\infty}^{t}\bigg{(}\bigcup_{k=\ell}^{\infty}E_{k}\cap B\bigg{)}\gg\mu(B).$
With the discussion at the beginning of Section 3, the proof of Theorem 1.4 is now complete.
4. Application to $\beta$-transformation
4.1. Definition and some basic properties
For $\beta>1$, the $\beta$-transformation $T_{\beta}:[0,1)\to[0,1)$ is defined by
$T_{\beta}x=\beta x\ (\textrm{mod}\ 1).$
For any $n\geq 1$ and $x\in[0,1)$, define $\epsilon_{n}(x,\beta)=\lfloor\beta T_{\beta}^{n-1}x\rfloor$, where $\lfloor x\rfloor$ denotes the integer part of $x$. Then, we can write
$x=\frac{\epsilon_{1}(x,\beta)}{\beta}+\frac{\epsilon_{2}(x,\beta)}{\beta^{2}}+\cdots+\frac{\epsilon_{n}(x,\beta)}{\beta^{n}}+\cdots,$
and we call the sequence
$\epsilon(x,\beta):=(\epsilon_{1}(x,\beta),\epsilon_{2}(x,\beta),\dots)$
the $\beta$-expansion of $x$. By the definition of $T_{\beta}$, it is clear that, for $n\geq 1$, $\epsilon_{n}(x,\beta)$ belongs to the alphabet $\{0,1,\dots,\lceil\beta-1\rceil\}$, where $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x$. When $\beta$ is not an integer, then not all sequences of $\{0,1,\dots,\lceil\beta-1\rceil\}^{\mathbb{N}}$ are the $\beta$-expansion of some $x\in[0,1)$. This leads to the notion of $\beta$-admissible sequence.
Definition 4.1.
A finite or an infinite sequence $(\epsilon_{1},\epsilon_{2},\dots)\in\{0,1,\dots,\lceil\beta-1\rceil\}^{\mathbb{N}}$ is said to be $\beta$-admissible if there exists an $x\in[0,1)$ such that the $\beta$-expansion of $x$ begins with $(\epsilon_{1},\epsilon_{2},\dots)$.
Denote by $\Sigma_{\beta}^{n}$ the collection of all admissible sequences of length $n$.
Definition 4.2.
For any $\bm{\epsilon}_{n}:=(\epsilon_{1},\dots,\epsilon_{n})\in\Sigma_{\beta}^{n}$, we call
$I_{n,\beta}(\bm{\epsilon}_{n}):=\{x\in[0,1):\epsilon_{k}(x,\beta)=\epsilon_{k},1\leq k\leq n\}$
an $n$th level cylinder.
Each cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$ can be viewed as a subinterval of $[0,1)$ consisting of all points whose first $n$ digits in their $\beta$-expansion coincide with the word $\bm{\epsilon}_{n}$. These cylinders form a natural partition of the interval $[0,1)$ at level $n$, and they shrink as $n$ increases. Clearly, for any $\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}$, the map $T_{\beta}^{n}$ is linear with slope $\beta^{n}$ when restricted to the cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$, and it sends $I_{n,\beta}(\bm{\epsilon}_{n})$ into $[0,1)$. If $\beta$ is not an integer, then the dynamical system $(T_{\beta},[0,1))$ is not a full shift, and thus $T_{\beta}^{n}|_{I_{n,\beta}(\bm{\epsilon}_{n})}$ may fail to be onto $[0,1)$. In other words, the length of $I_{n,\beta}(\bm{\epsilon}_{n})$ may be strictly less than $\beta^{-n}$, which complicates the analysis of the dynamical properties of $T_{\beta}$. In many cases, including the one considered here, it is more convenient to restrict attention to cylinders of maximal length, which motivates the definition of full cylinder.
Definition 4.3.
A cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$ or a sequence $\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}$ is called full if it has maximal length, that is, if
$|I_{n,\beta}(\bm{\epsilon}_{n})|=\beta^{-n}.$
In light of the definition of quasi-self-conformality, the collection $\mathcal{F}$ of sets required therein can naturally be taken to be the family of full cylinders. Moreover, in order to apply Theorem 1.4, it is necessary that the $\limsup$ set defined by full cylinders has full Lebesgue measure. Fortunately, this is indeed the case.
Lemma 4.4 ([38, Lemma 1 (1)]).
For any $N\geq 1$, we have
$\bigcup_{n=N}^{\infty}\bigcup_{\bm{\epsilon}_{n}\in\Lambda_{\beta}^{n}}I_{n,\beta}(\bm{\epsilon}_{n})=[0,1),$
where $\Lambda_{\beta}^{n}$ denotes the set of $n$th level full cylinders. In particular, the $\limsup$ set defined by all full cylinders has full Lebesgue measure.
The mass distribution principle stated in Proposition 2.1 requires estimating the measure of arbitrary balls in relation to their radii. However, after a detailed study of the distribution of full cylinders, Bugeaud and Wang [10, Proposition 1.3] showed that it suffices to consider balls that are themselves cylinders.
Proposition 4.5 (Modified mass distribution principle [10, Proposition 1.3]).
Let $E$ be a Borel measurable set in $[0,1]$ and $\lambda$ be a Borel measure with $\lambda(E)>0$. Assume that there exist a positive constant $c>0$ and an integer $n_{0}$ such that, for any $n\geq n_{0}$ the measure of any cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$ of order $n$ satisfies $\lambda(I_{n,\beta}(\bm{\epsilon}_{n}))\leq c|I_{n,\beta}(\bm{\epsilon}_{n})|^{s}$. Then, $\operatorname{\dim_{H}}E\geq s$.
4.2. Pressure function
Let $\phi:[0,1]\to\mathbb{R}$ be a continuous function. The pressure function for $\beta$-dynamical system associated to $\phi$ is defined by the limit
(4.1) $P(\phi,T_{\beta}):=\lim_{n\to\infty}\frac{1}{n}\log\sum_{\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}}e^{S_{n}\phi(y)},$
where for each admissible word $\bm{\epsilon}_{n}$, the point $y$ is any element in the corresponding cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$, and $S_{n}\phi(y)$ denotes the ergodic sum $\sum_{k=0}^{n-1}\phi(T_{\beta}^{k}y)$. The existence of the limit in (4.1) follows from the subadditivity:
$\log\sum_{\bm{\epsilon}_{n+m}\in\Sigma_{\beta}^{n+m}}e^{S_{n+m}\phi(y)}\leq\log\sum_{\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}}e^{S_{n}\phi(y)}+\log\sum_{\bm{\epsilon}_{m}\in\Sigma_{\beta}^{m}}e^{S_{m}\phi(T^{n}y)},$
and the limit does not depend on the choice of $y$ by the continuity of $g$.
It follows directly from the definition that the pressure function is continuous with respect to $\phi$.
Proposition 4.6.
Let $\phi$ and $\varphi$ be two continuous functions defined on $[0,1]$. Then,
$|P(\phi,T_{\beta})-P(\varphi,T_{\beta})|\leq\sup_{x\in[0,1]}|\phi(x)-\varphi(x)|.$
Consequently, if $\phi$ is positive, then there exists $0<s=s(\phi)<1$ such that
$P(-s(\phi+\log\beta),T_{\beta})=0.$
Guided by Theorem 1.4, it is necessary to choose a reference measure $\nu$ — generally singular with respect to the Lebesgue measure — to measure the size of the $\limsup$ sets. Such a measure is usually chosen as the Gibbs measure, whose existence is ensured by the following result.
Theorem 4.7 ([38, Theorems 13 and 16]).
Let $\phi:[0,1]\to\mathbb{R}$ be a Lipschitz continuous function. Then there exists a unique equilibrium state $\nu_{\phi}$ associated with $\phi$ such that the following properties hold:
  1. (1)
    The pressure satisfies the variational principle:
    $P(\phi,T_{\beta})=h_{\nu_{\phi}}+\int\phi\,d\nu_{\phi},$
    where $h_{\nu_{\phi}}$ is the measure-theoretic entropy of $\nu_{\phi}$ with respect to $T_{\beta}$.
  2. (2)
    For any cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$ of level $n$, the measure $\nu_{\phi}$ satisfies the Gibbs upper bound:
    $\nu_{\phi}(I_{n,\beta}(\bm{\epsilon}_{n}))\ll e^{S_{n}\phi(x)-nP(\phi,T_{\beta})},$
    where $x\in I_{n,\beta}(\bm{\epsilon}_{n})$ is arbitrary. Moreover, if the cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$ is full, then the Gibbs property holds in the sense:
    $\nu_{\phi}(I_{n,\beta}(\bm{\epsilon}_{n}))\asymp e^{S_{n}\phi(x)-nP(\phi,T_{\beta})}.$
It is important to note that in condition (3) of Theorem 1.4, one needs to compare the Hausdorff content $\mathcal{H}_{\infty}^{s}(E_{n})$ with $r_{n}^{\operatorname{\overline{\dim}_{H}}\nu}$. This comparison relies on understanding the Hausdorff dimension of the Gibbs measure $\nu_{\phi}$. The author believes that the following result has been established elsewhere; however, since no suitable reference could be found, we include the proof here.
Lemma 4.8.
Let $\phi:[0,1]\to\mathbb{R}$ be a Lipschitz continuous function and $\nu_{\phi}$ be the associated equilibrium state. Then,
$\operatorname{\dim_{H}}\nu_{\phi}=\frac{h_{\nu_{\phi}}}{\log\beta}.$
Proof.
By Birkhoff’s ergodic theorem, for $\nu_{\phi}$ almost all $x$,
(4.2) $\lim_{n\to\infty}\frac{1}{n}S_{n}\phi(x)=\int\phi\,\mathrm{d}\nu_{\phi}.$
For any $n\geq 1$, denote by $I_{n,\beta}(x)$ the cylinder of level $n$ that contains $x$. For any $x$ for which (4.2) holds, by the definition of local dimension,
$\begin{split}\underline{D}(\nu_{\phi},x)&=\liminf_{r\to 0}\frac{\log\nu_{\phi}(B(x,r))}{\log r}\leq\liminf_{n\to\infty}\frac{\log\nu_{\phi}(I_{n,\beta}(x))}{\log|I_{n,\beta}(x)|}\\ &\leq\liminf_{\begin{subarray}{c}n\to\infty\\ I_{n,\beta}(x)\text{ if full}\end{subarray}}\frac{\log\nu_{\phi}(I_{n,\beta}(x))}{\log|I_{n,\beta}(x)|},\end{split}$
where, in the last inequality, we use the fact that for any $x\in[0,1)$, there exists infinitely many $n$ such that $I_{n,\beta}(x)$ is full (see Lemma 4.4). For any full cylinder $I_{n,\beta}(x)$, by the Gibbs property,
$\nu_{\phi}(I_{n,\beta}(x))\asymp e^{S_{n}\phi(x)-nP(\phi,T_{\beta})}=e^{n(\int\phi\,\mathrm{d}\nu_{\phi}-P(\phi,T_{\beta})+o(1))},$
where $o(1)\to 0$ as $n\to\infty$. Therefore,
$\begin{split}\underline{D}(\nu_{\phi},x)&\leq\liminf_{\begin{subarray}{c}n\to\infty\\ I_{n,\beta}(x)\text{ if full}\end{subarray}}\frac{\int\phi\,\mathrm{d}\nu_{\phi}-P(\phi,T_{\beta})+o(1)}{-\log\beta}=\frac{P(\phi,T_{\beta})-\int\phi\,\mathrm{d}\nu_{\phi}}{\log\beta}\\ &=\frac{h_{\nu_{\phi}}}{\log\beta},\end{split}$
where the equality follows from the variational principle. Since this holds for $\nu_{\phi}$ almost all $x$, we have
(4.3) $\operatorname{\overline{\dim}_{H}}\nu_{\phi}\leq\frac{h_{\nu_{\phi}}}{\log\beta}.$
Next, we prove the reverse inequality. For any set $E$ with positive $\nu_{\phi}$-measure, we claim that
(4.4) $\operatorname{\dim_{H}}E\geq\frac{h_{\nu_{\phi}}}{\log\beta}.$
It follows from the definition that
$\operatorname{\underline{\dim}_{H}}\nu_{\phi}\geq\frac{h_{\nu_{\phi}}}{\log\beta},$
which together with (4.3) concludes the proof.
Fix a Borel set $E$ with $\nu_{\phi}(E)>0$. Let $\varepsilon>0$. Then, there exists an integer $N=N(\varepsilon)$ such that the set $A_{N}$ of $x$ for which
$\bigg{|}\frac{1}{n}S_{n}\phi(x)-\int\phi\,\mathrm{d}\nu_{\phi}\bigg{|}<\varepsilon,\quad\text{for all $n\geq N$}$
has $\nu_{\phi}$-measure larger than $1-\nu_{\phi}(E)$. Obviously, $A_{N}\cap E$ has positive $\nu_{\phi}$-measure.
Let $\lambda=\nu_{\phi}|_{A_{N}\cap E}$. For any $n\geq N$ and any cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$, if
$\bigg{|}\frac{1}{n}S_{n}\phi(x)-\int\phi\,\mathrm{d}\nu_{\phi}\bigg{|}\geq\varepsilon$
for all $x\in I_{n,\beta}(\bm{\epsilon}_{n})$. Then,
$\lambda(I_{n,\beta}(\bm{\epsilon}_{n}))=\nu_{\phi}(A_{N}\cap E\cap I_{n,\beta}(\bm{\epsilon}_{n}))=0.$
Therefore, for any cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$,
$\lambda(A_{N}\cap E\cap I_{n}(\bm{\epsilon}_{n}))>0$
(4.5) $\Longrightarrow\quad$ $\bigg{|}\frac{1}{n}S_{n}\phi(x)-\int\phi\,\mathrm{d}\nu_{\phi}\bigg{|}<\varepsilon\text{ for some $x\in I_{n,\beta}(\bm{\epsilon}_{n})$}.$
We stress that the reverse implication may not be true. Let $I_{n,\beta}(\bm{\epsilon}_{n})$ be a cylinder with positive $\nu_{\phi}$ measure and let $x=x(\bm{\epsilon}_{n})$ be such that (4.5) holds. Let $m\geq n$ be the unique integer satisfy
$\beta^{-m-1}<|I_{n,\beta}(\bm{\epsilon}_{n})|\leq\beta^{-m}.$
It follows that $I_{n,\beta}(\bm{\epsilon}_{n})=I_{m,\beta}(\bm{\epsilon}_{n},0^{m-n})$, where $0^{m-n}$ denotes the word consisting of mnm-n consecutive zeros. Since $x\in I_{n,\beta}(\bm{\epsilon}_{n})=I_{m,\beta}(\bm{\epsilon}_{n},0^{m-n})$, it follows from Theorem 4.7 that
$\begin{split}\lambda(I_{n,\beta}(\bm{\epsilon}_{n}))&\leq\nu_{\phi}(I_{m,\beta}(\bm{\epsilon}_{n},0^{m-n}))\ll e^{S_{m}\phi(x)-mP(\phi,T_{\beta})}\\ &\leq e^{m\int\phi\,\mathrm{d}\nu_{\phi}+m\varepsilon-mP(\phi,T_{\beta})}=\beta^{-m(P(\phi,T_{\beta})-\int\phi\,\mathrm{d}\nu_{\phi}-\varepsilon)/\log\beta}\\ &\asymp|I_{n,\beta}(\bm{\epsilon}_{n})|^{(P(\phi,T_{\beta})-\int\phi\,\mathrm{d}\nu_{\phi}-\varepsilon)/\log\beta}.\end{split}$
By Proposition 4.5,
$\operatorname{\dim_{H}}(A_{N}\cap E)\geq\frac{P(\phi,T_{\beta})-\int\phi\,\mathrm{d}\nu_{\phi}-\varepsilon}{\log\beta}=\frac{h_{\nu_{\phi}}-\varepsilon}{\log\beta}.$
By the arbitrariness of $\varepsilon$, the claim (4.4) follows immediately. ∎
4.3. Application to shrinking target problems
Recall that $\{h_{n}\}$ is a sequence of Lipschitz functions with uniformly bounded Lipschitz constants, and that
$W(T_{\beta},f,\{h_{n}\})=\{x\in[0,1):|T_{\beta}^{n}x-h_{n}(x)|<e^{-S_{n}f(x)}\text{ for i.m. $n$}\},$
where $f:[0,1]\to\mathbb{R}$ is a positive continuous function. In this section, we will prove that
(4.6) $W(T_{\beta},f,\{h_{n}\})\in\mathscr{G}^{s}([0,1]),$
where $s$ satisfies $P(-s(f+\log\beta),T_{\beta})=0$.
Proof of (4.6).
Note that Lipschitz functions are dense in $C^{0}([0,1])$. Approximating $f$ from above and using the continuity of pressure function, we can assume that $f$ is Lipschitz. Let $\nu_{s}$ be the Gibbs measure associated to the Lipschitz function $s(f+\log\beta)$. The Gibbs property of $\nu_{s}$ (see Theorem 4.7) ensures that $\nu_{s}$ is quasi-self-conformal with respect to the collection of full cylinders. Moreover, by Lemma 4.4, the $\limsup$ set defined by the collection of full cylinders has full Lebesgue. Therefore, Theorem 1.4 can be applied to $\nu_{s}$.
Birkhoff’s ergodic theorem gives that for $\nu_{s}$ almost all $x$,
$\lim_{n\to\infty}\frac{1}{n}S_{n}f(x)=\int f\,\mathrm{d}\nu_{s}.$
Since for any $x\in[0,1)$, there exist infinitely many $n$ such that $I_{n,\beta}(x)$ is full, it is not difficult to verify that the set
$\bigcap_{N=1}^{\infty}\bigcup_{n=N}^{\infty}\bigcup_{\bm{\epsilon}_{n}\in\Lambda_{\beta}^{n}(\nu_{s},\varepsilon)}I_{n,\beta}(\bm{\epsilon}_{n})$
is of full $\nu_{s}$-measure, where recall that $\Lambda_{\beta}^{n}$ is the set of $n$th level full sequences, and
$\Lambda_{\beta}^{n}(\nu_{s},\varepsilon):=\bigg{\{}\bm{\epsilon}_{n}\in\Lambda_{\beta}^{n}:\bigg{|}\frac{1}{n}S_{n}f(x)-\int f\,\mathrm{d}\nu_{s}\bigg{|}<\varepsilon\text{ for all $x\in I_{n,\beta}(\bm{\epsilon}_{n})$}\bigg{\}}.$
Then,
$\begin{split}W(T_{\beta},f,\{h_{n}\})\supset\bigcap_{N=1}^{\infty}\bigcup_{n=N}^{\infty}\bigcup_{\bm{\epsilon}_{n}\in\Lambda_{\beta}^{n}(\nu_{s},\varepsilon)}I_{n,\beta}(\bm{\epsilon}_{n})\cap E_{n}(T_{\beta},f,h_{n}),\end{split}$
where $E_{n}(T_{\beta},f,h_{n})=\{x\in[0,1):|T_{\beta}^{n}x-h_{n}(x)|<e^{-S_{n}f(x)}\}$. We need the following lemma to estimate the size of $I_{n,\beta}(\bm{\epsilon}_{n})\cap E_{n}(T_{\beta},f,h_{n})$.
Lemma 4.9 ([42]).
Let $h$ be a Lipschitz function with Lipschitz constant $L\geq 0$. Let $0<r<1$. For any $n$ with $L<\beta^{n}$ and any sequence $\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}$, the set
$\{x\in I_{n,\beta}(\bm{\epsilon}_{n}):|T_{\beta}^{n}x-h(x)|<r\}$
is contained in a ball of radius $2r\beta^{-n}$. Moreover, if $\bm{\epsilon}_{n}$ is full, then it contains a ball of radius $r\beta^{-n}/2$.
By Lipschitz continuity, for any $x,y\in I_{n,\beta}(\bm{\epsilon}_{n})$,
$|S_{n}f(x)-S_{n}f(y)|\ll 1\quad\Longrightarrow\quad e^{-S_{n}f(x)}\asymp e^{-S_{n}f(y)}.$
Therefore, since $\bm{\epsilon}_{n}\in\Lambda_{\beta}^{n}(\nu_{s},\varepsilon)$ is full, we can apply the above lemma to conclude that $I_{n,\beta}(\bm{\epsilon}_{n})\cap E_{n}(T_{\beta},f,h_{n})$ contains an interval of length
$\asymp\beta^{-n}e^{-S_{n}f(x)}\geq\beta^{-n}e^{-n\int f\,\mathrm{d}\nu_{s}-n\varepsilon}=\beta^{-n-n(\int f\,\mathrm{d}\nu_{s}+\varepsilon)/\log\beta}.$
By the variational principle and note that $P(-s(f+\log\beta),T_{\beta})=0$, we have
$0=h_{\nu_{s}}-s\bigg{(}\int f\,\mathrm{d}\nu_{s}+\log\beta\bigg{)}\quad\Longrightarrow\quad s=\frac{h_{\nu_{s}}}{\int f\,\mathrm{d}\nu_{s}+\log\beta}.$
Let $K$ be a sufficiently large integer, independent of $\varepsilon$, such that the following inequality holds:
$\bigg{(}1+\frac{\int f\,\mathrm{d}\nu_{s}+\varepsilon}{\log\beta}\bigg{)}\bigg{(}\frac{h_{\nu_{s}}}{\int f\,\mathrm{d}\nu_{s}+\log\beta}-K\varepsilon\bigg{)}\leq\frac{h_{\nu_{s}}}{\log\beta}.$
Then, by Lemma 4.8,
$\begin{split}\mathcal{H}_{\infty}^{s-K\varepsilon}\big{(}I_{n,\beta}(\bm{\epsilon}_{n})\cap E_{n}(T_{\beta},f,h_{n})\big{)}&\gg\beta^{-n(1+(\int f\,\mathrm{d}\nu_{s}+\varepsilon)/\log\beta)(s-K\varepsilon)}\\ &\geq\beta^{-nh_{\nu_{s}}/\log\beta}=|I_{n,\beta}(\bm{\epsilon}_{n})|^{-\operatorname{\dim_{H}}\nu_{s}}.\end{split}$
Therefore, by Theorem 1.4, we have
$W(T_{\beta},f,\{h_{n}\})\in\mathscr{G}^{s-K\varepsilon}([0,1]).$
Since $\varepsilon$ is arbitrary and $K$ does not depend on $\varepsilon$,
$W(T_{\beta},f,\{h_{n}\})\in\mathscr{G}^{s}([0,1]).\qed$
5. Applications to Gauss map
5.1. Definition and some basic properties
The Gauss map $G:[0,1)\to[0,1)$ is defined by
$G(0):=0\mathrm{\quad and\quad}G(x)=\frac{1}{x}\ (\textrm{mod}\ 1)\quad\text{for $x\in(0,1)$}.$
It is well-known that every irrational $x\in(0,1)$ can be written uniquely as an infinite expansion of the form
(5.1) $x=\frac{1}{a_{1}(x)+\cfrac{1}{a_{2}(x)+\cfrac{1}{a_{3}(x)+\cdots}}}=:[a_{1}(x),a_{2}(x),a_{3}(x),\dots],$
where $a_{1}(x)=\lfloor 1/x\rfloor$ and $a_{n}(x)=a_{1}(G^{n-1}x)$ for $n\geq 2$ are called the partial quotients of $x$. The $n$th truncation $a_{1}(x),\dots,a_{n}(x)]$, denoted by $p_{n}(x)/q_{n}(x)$ is called the $n$th convengent of $x$. With the convention
$p_{-1}=1,\quad q_{-1}=0,\quad p_{0}=0\quad\text{and}\quad q_{0}=1,$
the convergents $\{p_{n}/q_{n}\}=\{p_{n}(x)/q_{n}(x)\}$ of $x$ can be generated by the recursive formulae:
(5.2) $p_{n}=a_{n}p_{n-1}+p_{n-2}\quad\text{and}\quad q_{n}=a_{n}q_{n-1}+q_{n-2}\qquad\text{for $n\geq 1$}.$
These expressions show that both $p_{n}$ and $q_{n}$ are completely determined by the initial segment $\bm{a}_{n}:=(a_{1},\dots,a_{n})\in\mathbb{N}^{n}$ of partial quotients. We therefore write
$p(\bm{a}_{n})=p_{n}\quad\text{and}\quad q(\bm{a}_{n})=q_{n}.$
Definition 5.1.
For any $\bm{a}_{n}\in\mathbb{N}^{n}$, we call
$I_{n}(\bm{a}_{n}):=\{x\in[0,1):a_{k}(x)=a_{k},1\leq k\leq n\}$
an $n$th level cylinder.
Geometrically, these cylinders form a nested partition of the unit interval, refining as $n$ increases. The length of each cylinder decays exponentially with $n$ and can be precisely estimated in terms of the denominators $q_{n}$ of the convergents:
Lemma 5.2 ([27, 24]).
Let $\bm{a}_{n}\in\mathbb{N}^{n}$. Then the corresponding $n$th level cylinder satisfies the bounds
$q(\bm{a}_{n})^{-2}/2<|I_{n}(\bm{a}_{n})|\leq q(\bm{a}_{n})^{-2},$
and moreover,
$|I_{n}(\bm{a}_{n})|\asymp e^{-S_{n}\log|G^{\prime}|(x)}$
where $x$ belongs to the interior of $I_{n}(\bm{a}_{n})$.
The next proposition describes the positions of cylinders of level $n+1$ inside the $n$th level cylinder.
Proposition 5.3 ([27]).
Let $I_{n}(\bm{a}_{n})$ be an $n$th level cylinder, which is partitioned into sub-cylinders $\{I_{n+1}(\bm{a}_{n},a_{n+1}):a_{n+1}\in\mathbb{N}\}$. When $n$ is odd, these sub-cylinders are positioned from left to right, as $a_{n+1}$ increases from $1$ to $\infty$; when $n$ is even, they are positioned from right to left.
5.2. Pressure function
Let $\phi:[0,1]\to\mathbb{R}$ be a Lipschitz continuous function. The pressure function for Gauss map associated to $\phi$ is defined as
(5.3) $P(\phi,G):=\lim_{n\to\infty}\frac{1}{n}\log\sum_{\bm{a}_{n}\in\mathbb{N}^{n}}e^{S_{n}\phi(y)},$
where $y\in I_{n}(\bm{a}_{n})$. The proof of the existence of limit in (5.3) can be found in [32, Proposition 2.4]. Compared with the $\beta$-transformation, one major difference is that there are infinitely many $n$th level cylinders. As a result, the summation in (5.3) may be infinite, and hence the pressure function may fail to be continuous with respect to $\phi$. For this reason, instead of providing a comprehensive but technically involved description of the pressure function for the Gauss map, we merely summarize part of the results from [40, 34, 18] and refer the reader to these references for further details.
Theorem 5.4 ([40, 34, 18]).
Let $f:[0,1]\to\mathbb{R}$ be a positive Lipschitz function. Then the function $t\mapsto P(-t(f+\log|G^{\prime}|),G)$ is continuous on $(1/2,+\infty)$. Moreover, there exists $t=t(f)\in(1/2,1)$ such that
$P(-t(f+\log|G^{\prime}|),G)=0.$
For the function $t(f+\log|G^{\prime}|)$, there exists a unique equilibrium state $\nu_{t}$ satisfying
$0=h_{\nu_{t}}-t\int(f+\log|G^{\prime}|)\,\mathrm{d}\nu_{t},$
and such that, for any $\bm{a}_{n}\in\mathbb{N}^{n}$ and any $x\in I_{n}(\bm{a}_{n})$,
$\nu_{t}(I_{n}(\bm{a}_{n}))\asymp e^{-ntS_{n}(f+\log|G^{\prime}|)(x)}.$
Sketch of the proof.
Note that the continued fraction dynamical system can be viewed as an iterated function system:
$S=\bigg{\{}\phi_{i}(x)=\frac{1}{i+x}:i\in\mathbb{N}\bigg{\}}.$
It then follows from [34, Proposition 3.3] that the pressure function $P(-t\log|G^{\prime}|,G)$ is continuous on $(1/2,\infty)$, since by [40, Lemma 2.6], we have
$\lim_{u\to 1/2^{+}}P(-u\log|G^{\prime}|,G)=\infty.$
Now, observe that
$\begin{split}P(-t\log|G^{\prime}|,G)-t\max_{x\in[0,1]}f(x)&\leq P(-t(f+\log|G^{\prime}|),G)\\ &\leq P(-t\log|G^{\prime}|,G)-t\min_{x\in[0,1]}f(x).\end{split}$
Since $f$ is positive and $P(-\log|G^{\prime}|,G)=0$, and noting that the pressure function is strictly decreasing in $t$, there must exist a unique $t\in(1/2,1)$ such that
$P\left(-t(f+\log|G^{\prime}|),G\right)=0.$
Finally, the Lipschitz continuity of $f$ ensures the existence and uniqueness of the equilibrium state associated with the potential $t(f+\log|G^{\prime}|)$; see [18, Theorem 2.16]. In addition, this equilibrium state satisfies the corresponding Gibbs property; see [18, (2.16’)]. ∎
It follows from ergodic theorem and Lemma [17, Lemma 2.12 (b)] that for the Gibbs measure $\nu_{t}$ given in Theorem 5.4,
(5.4) $\operatorname{\dim_{H}}\nu_{t}=\frac{h_{\nu_{t}}}{\int\log|T^{\prime}|\,\mathrm{d}\nu_{t}}.$
5.3. Applications
Note that for any $\bm{a}_{n}\in\mathbb{N}^{n}$, it is known (see, e.g., [26, Lemma 2.5]) that
$G^{n}|_{I_{n}(\bm{a}_{n})}=[0,1)\mathrm{\quad and\quad}|(G^{n})^{\prime}(x)|\asymp q(\bm{a}_{n})^{2}.$
The Gibbs property further implies that $\nu_{t}$ is quasi-self-conformal with respect to the collection of all cylinders. By a suitable arrangement, it is easy to verify that the $\limsup$ set defined by the collection of cylinders has full Lebesgue measure. This key observation enables us to apply Theorem 1.4 to the measure $\nu_{t}$. In light of Theorem 5.4 and the dimension formula (5.4), we conclude — by arguments similar to those used in the proof of the large intersection property of $W(T_{\beta},f,\{h_{n}\})$ (see §4.3)— that for any continuous positive function $f:[0,1]\to\mathbb{R}$,
$W(G,f,\{h_{n}\})\in\mathscr{G}^{t}([0,1]),$
where $t$ solves the pressure equation $P(-t(f+\log|G^{\prime}|),G)=0$.
We now turn to another class of sets defined in terms of growth conditions on blocks of consecutive partial quotients. Recall that for any integer $m\geq 1$ and real number $B>1$, we define
$F_{m}(B):=\{x\in[0,1):a_{n+1}(x)\cdots a_{n+m}(x)\geq B^{n}\text{ for i.m. $n$}\}.$
Our goal in the remainder of this subsection is to prove that
(5.5) $F_{m}(B)\in\mathscr{G}^{u}([0,1]),$
for some $u\in(1/2,1)$ satisfying
(5.6) $P(-u\log|G^{\prime}|-g_{m}(u)\log B,G)=0,$
where the function $g_{m}(u)$ is given by
$g_{m}(u)=\frac{u^{m}(2u-1)}{u^{m}-(1-u)^{m}}.$
The existence of $u$ satisfying (5.6) is ensured by the following lemma, which follows from standard properties of the pressure function.
Lemma 5.5.
Let $m\geq 1$ be an integer and $B>1$. There exists $1/2<u<1$ such that
$P(-u\log|G^{\prime}|-g_{m}(u)\log B,G)=0.$
Proof.
By the definition of the pressure function, we can write
$P(-u\log|G^{\prime}|-g_{m}(u)\log B,G)=P(-u\log|G^{\prime}|,G)-g_{m}(u)\log B.$
Let us now consider the two functions appearing on the right-hand-side. On the one hand, as shown in the proof of [40, Lemma 2.6],
$\lim_{u\to 1/2^{+}}P(-u\log|G^{\prime}|,G)=\infty,\quad\text{while}\quad P(-\log|G^{\prime}|,G)=0.$
On the other hand, note that $g_{m}(u)$ is continuous on $(1/2,1)$, and satisfies
$g_{m}(1/2)\log B=0,\quad\text{and}\quad g_{m}(1)\log B=\log B>0.$
Combining this with the continuity of both functions, it follows that the function $u\mapsto P(-u\log|G^{\prime}|,G)-g_{m}(u)\log B$ is continuous and takes values $\infty$ near $u=1/2$ and negative near $u=1$. By the intermediate value theorem, there exists some $u\in(1/2,1)$ such that the equation stated in the lemma equals zero. ∎
Let $u$ be as in Theorem 1.6. Denote by $\nu_{u}$ the Gibbs measure associated to $u\log|G^{\prime}|-g_{m}(u)\log B$. By the dimension formula (5.4), we have
$\operatorname{\dim_{H}}\nu_{u}=\frac{h_{\nu_{u}}}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}.$
By Birkhoff’s ergodic theorem, for $\nu_{u}$ almost all $x$,
$\lim_{n\to\infty}\frac{1}{n}S_{n}\log|G^{\prime}|(x)=\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}.$
For any $\varepsilon>0$, it is not difficult to verify that the set
$\bigcap_{N=1}^{\infty}\bigcup_{n=N}^{\infty}\bigcup_{\bm{a}_{n}\in\Gamma_{n}(\nu_{u},\varepsilon)}I_{n}(\bm{a}_{n})$
is of full $\nu_{u}$-measure, where
$\Gamma_{n}(\nu_{u},\varepsilon):=\bigg{\{}\bm{a}_{n}\in\mathbb{N}^{n}:\bigg{|}\frac{1}{n}S_{n}\log|G^{\prime}|(x)-\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}\bigg{|}<\varepsilon\text{ for all $x\in I_{n}(\bm{a}_{n})$}\bigg{\}}.$
Consequently, we obtain the inclusion
$\begin{split}F_{m}(B)\supset\bigcap_{N=1}^{\infty}\bigcup_{n=N}^{\infty}\bigcup_{\bm{a}_{n}\in\Gamma_{n}(\nu_{u},\varepsilon)}\{x\in I_{n}(\bm{a}_{n}):a_{n+1}(x)\cdots a_{n+m}(x)\geq B^{n}\}.\end{split}$
To analyze the growth condition on the partial quotients, define the sequence
$\alpha_{i}=B^{g_{m}(u)(1-u)^{i-1}u^{-i}}\quad\text{for $1\leq i\leq m-1$},$
and set
$\alpha_{m}=\frac{B}{\alpha_{1}\cdots\alpha_{m-1}}.$
It can be deduced from the expression of $g_{m}(u)$ that the following equalities hold:
(5.7) $\alpha_{1}^{u}=\alpha_{1}^{2u-1}\alpha_{2}^{u}=\cdots=(\alpha_{1}\cdots\alpha_{m-1})^{2u-1}\alpha_{m}^{u}=B^{g_{m}(u)}.$
From now on, fix $\bm{a}_{n}\in\Gamma_{n}(\nu_{u},\varepsilon)$. We construct an open set inside $\{x\in I_{n}(\bm{a}_{n}):a_{n+1}(x)\cdots a_{n+m}(x)\geq B^{n}\}$ as follows:
(5.8) $A:=\{x\in I_{n}(\bm{a}_{n}):\alpha_{i}^{n}\leq a_{n+i}(x)\leq 2\alpha_{i}^{n}\text{ and $a_{n+i}(x)$ is even for $1\leq i\leq m$}\}.$
Here, we require that $a_{n+i}(x)$ is even to ensure that cylinders of level $n+m$ contained in $A$ are well-separated, in the sense described below.
Lemma 5.6.
Let $I_{n+m}(\bm{a}_{n},a_{n+1},\dots,a_{n+m})$ and $I_{n+m}(\bm{a}_{n},a_{n+1}^{\prime},\dots,a_{n+m}^{\prime})$ be two distinct cylinders contained in $A$. Let $1\leq k\leq m$ be the smallest integer for which $a_{n+k}\neq a_{n+k}^{\prime}$. Then, the distance between these two cylinders is at least
$\frac{1}{32q(\bm{a}_{n},a_{n+1},\dots,a_{n+k})^{2}}.$
Proof.
By the distribution properties of cylinders (see Proposition 5.3) and the fact that by definition both $a_{n+k}$ and $a_{n+k}^{\prime}$ are even integers, there exists a cylinder
$I_{n+k}(\bm{a}_{n},a_{n+1},\dots,a_{n+k-1},a_{n+k}^{\prime\prime})$
with either $a_{n+k}<a_{n+k}^{\prime\prime}<a_{n+k}^{\prime}$ or $a_{n+k}^{\prime}<a_{n+k}^{\prime\prime}<a_{n+k}$, which lies between the two cylinders stated in the lemma. Therefore, by Lemma 5.2, (5.2) and (5.8), they are separated by a distance
$\begin{split}|I_{n+k}(\bm{a}_{n},a_{n+1},\dots,a_{n+k-1},a_{n+k}^{\prime\prime})|&\geq\frac{1}{2q(\bm{a},a_{n+1},\dots,a_{n+k-1},a_{n+k}^{\prime\prime})^{2}}\\ &\geq\frac{1}{32q(\bm{a},a_{n+1},\dots,a_{n+k-1},a_{n+k})^{2}},\end{split}$
which provides the claimed lower bound on the distance between the two cylinders. ∎
Define a probability measure $\lambda$ supported on $A$ by
(5.9) $\lambda=\frac{1}{\#A}\sum_{I_{n+m}(\bm{a}_{n+m})\subset A}\frac{\mathcal{L}|_{I_{n+m}(\bm{a}_{n+m})}}{\mathcal{L}(I_{n+m}(\bm{a}_{n+m}))},$
where $\mathcal{L}$ denotes the one-dimensional Lebesgue measure.
Lemma 5.7.
Let $\lambda$ be as above. For any $x\in A$ and $r>0$, we have
$\lambda(B(x,r))\ll r^{u-K\varepsilon}q(\bm{a}_{n})^{2(u-K\varepsilon)}B^{ng_{m}(u)},$
where $K$ is a sufficiently large integer, independent of $\varepsilon$, such that
$u-K\varepsilon+\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}-\varepsilon}\leq u+\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}.$
Proof.
Without loss of generality, assume that $x\in I_{n+m}:=I_{n+m}(\bm{a}_{n},a_{n+1},\dots,a_{n+m})\subset A$. Obviously, if $r$ is relatively large, specifically
$r\geq\frac{1}{32q(\bm{a}_{n},a_{n+1}\dots,a_{n+m})^{2}}\geq\frac{|I_{n}(\bm{a}_{n})|}{32},$
then trivially,
$\lambda(B(x,r))\leq 1\ll\frac{r^{u-K\varepsilon}}{|I_{n}(\bm{a}_{n})|^{u-K\varepsilon}}\ll r^{u-K\varepsilon}q(\bm{a}_{n})^{2(u-K\varepsilon)}B^{ng_{m}(u)}.$
Hence, it is sufficient to focus on the case $r<1/(32\,q(\bm{a}_{n},a_{n+1},\dots,a_{n+m})^{2})$. By Lemma 5.6, the cylinders in $A$ are well-separated, allowing us to focus on two distinct cases.
Case 1: Suppose there exists some $1\leq k\leq m$ such that
$\frac{1}{32\,q(\bm{a}_{n},a_{n+1},\dots,a_{n+k})^{2}}\leq r<\frac{1}{32\,q(\bm{a}_{n},a_{n+1},\dots,a_{n+k-1})^{2}}.$
By Lemma 5.6, the ball $B(x,r)$ only intersects one cylinder of level n+k1n+k-1, namely $I_{n+k-1}(\bm{a}_{n},a_{n+1}\dots,a_{n+k-1})$, contained in $A$, but may intersect multiple cylinders of level $n+k$. Define
$\begin{split}\Delta(x;k)=\{a_{n+k}\in[\alpha_{k}^{n},2\alpha_{k}^{n}]:&\text{ $a_{n+k}$ is even and }\\ &I_{n+k}(\bm{a}_{n},a_{n+1}\dots,a_{n+k-1},a_{n+k})\cap B(x,r)\neq\emptyset\}.\end{split}$
To estimate $\mu(B(x,r))$, it is essential to bound $\Delta(x;k)$ from above. Two natural upper bounds arise:
  1. (a)
    From the definition,
    (5.10) $\Delta(x;k)\ll\alpha_{k}^{n}.$
  2. (b)
    From the well-separation property (Lemma 5.6), cylinders of level $n+k$ in $A$ are spaced by at least $1/(32\,q(\bm{a}_{n},a_{n+1},\dots,a_{n+k})^{2})$. Thus,
    (5.11) $\Delta(x;k)\ll r\,q(\bm{a}_{n},a_{n+1},\dots,a_{n+k})^{2}\asymp r\,q(\bm{a}_{n})^{2}a_{n+1}^{2}\cdots a_{n+k}^{2}.$
Combining (5.10) and (5.11) and using the inequality $\max\{a,b\}\leq a^{1-u}b^{u}$, we get
$\begin{split}\#\Delta(x;k)&\ll\max\{\alpha_{k}^{n},rq(\bm{a}_{n})^{2}a_{n+1}^{2}\cdots a_{n+k}^{2}\}\\ &\ll\alpha_{k}^{n(1-u)}\cdot\left(rq(\bm{a}_{n})^{2}a_{n+1}^{2}\cdots a_{n+k}^{2}\right)^{u}\\ &\ll r^{u}q(\bm{a}_{n})^{2u}\alpha_{1}^{2nu}\cdots\alpha_{k-1}^{2nu}\alpha_{k}^{n(1+u)},\end{split}$
where we use $a_{n+i}\asymp\alpha_{i}^{n}$ for $1\leq i\leq m$ in the last inequality. Therefore, by the definition of $\lambda$ and the equalities presented in (5.7),
$\begin{split}\lambda(B(x,r))&\ll\#\Delta(x;k)\cdot\frac{1}{\alpha_{1}^{n}\cdots\alpha_{k}^{n}}\\ &\ll r^{u}q(\bm{a}_{n})^{2u}\alpha_{1}^{n(2u-1)}\cdots\alpha_{k-1}^{n(2u-1)}\alpha_{k}^{nu}\\ &=r^{u}q(\bm{a}_{n})^{2u}B^{ng_{m}(u)}\ll r^{u-K\varepsilon}q(\bm{a}_{n})^{2(u-K\varepsilon)}B^{ng_{m}(u)}.\end{split}$
Case 2: If
$r\leq\frac{1}{32\,q(\bm{a}_{n},a_{n+1},\dots,a_{n+m})^{2}},$
then $B(x,r)$ intersects only one cylinder of level $n+m$, namely $I_{n+m}$, contained in $A$. It follows that
$\begin{split}\frac{\mathcal{L}|_{I_{n+m}}(B(x,r))}{\mathcal{L}(I_{n+m})}&\leq\frac{2r}{\mathcal{L}(I_{n+m})}\ll rq(\bm{a}_{n},a_{n+1},\dots,a_{n+m})^{2}\\ &\ll r^{u}q(\bm{a}_{n},a_{n+1},\dots,a_{n+m})^{2u}\asymp r^{u}q(\bm{a}_{n})^{2u}\alpha_{1}^{2nu}\cdots\alpha_{m}^{2nu}\\ &\leq r^{u}q(\bm{a}_{n})^{2u}\alpha_{1}^{2nu}\cdots\alpha_{m-1}^{2nu}\alpha_{m}^{n(u+1)},\end{split}$
where we use $1/2<u<1$ in the last inequality. Again, by the definition of $\lambda$ and the equalities presented in (5.7),
$\begin{split}\lambda(B(x,r))&=\frac{1}{\#A}\cdot\frac{\mathcal{L}|_{I_{n+m}}(B(x,r))}{\mathcal{L}(I_{n+m})}\ll r^{u}q(\bm{a}_{n})^{2u}\alpha_{1}^{n(2u-1)}\cdots\alpha_{m-1}^{n(2u-1)}\alpha_{m}^{nu}\\ &=r^{u}q(\bm{a}_{n})^{2u}B^{ng_{m}(u)}\ll r^{u-K\varepsilon}q(\bm{a}_{n})^{2(u-K\varepsilon)}B^{ng_{m}(u)}.\qed\end{split}$
We are now in a position to prove Theorem 1.6 (see also (5.5)), using the measure $\lambda$ constructed earlier and the mass distribution principle.
Proof of Theorem 1.6.
Let $\bm{a}_{n}\in\Gamma_{n}(\nu_{u},\varepsilon)$. By Lemma 5.7 and the mass distribution principle, we obtain the following lower bound for the $(u-K\varepsilon)$-Hausdorff content of the set $A$:
(5.12) $\mathcal{H}_{\infty}^{u-K\varepsilon}(A)\gg q(\bm{a}_{n})^{-2(u-K\varepsilon)}B^{-ng_{m}(u)}\asymp e^{-(u-K\varepsilon)S_{n}\log|G^{\prime}|(x)-ng_{m}(u)\log B},$
where $x\in I_{n}(\bm{a}_{n})$ is any point in the cylinder. To proceed, we analyze the exponent on the right-hand-side of (5.12). By the definition of $\Gamma_{n}(\nu_{u},\varepsilon)$, we know that for any $x\in I_{n}(\bm{a}_{n})$,
$\bigg{|}\frac{1}{n}S_{n}\log|G^{\prime}|(x)-\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}\bigg{|}<\varepsilon\quad\Longrightarrow\quad S_{n}\log|G^{\prime}|(x)\geq n\bigg{(}\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}-\varepsilon\bigg{)}.$
Substituting this into the exponent, we obtain:
$(u-K\varepsilon)S_{n}\log|G^{\prime}|(x)+ng_{m}(u)\log B$
$=$ $S_{n}\log|G^{\prime}|(x)\bigg{(}u-K\varepsilon+\frac{ng_{m}(u)\log B}{S_{n}\log|G^{\prime}|(x)}\bigg{)}$
$\leq$ $S_{n}\log|G^{\prime}|(x)\bigg{(}u-K\varepsilon+\frac{ng_{m}(u)\log B}{n(\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}-\varepsilon)}\bigg{)}$
(5.13) $\leq$ $S_{n}\log|G^{\prime}|(x)\bigg{(}u+\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}\bigg{)},$
where the final inequality follows from the choice of the constant $K$. By the variational principle, we have
$\begin{split}&0=h_{\nu_{u}}-\bigg{(}u\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}+g_{m}(u)\log B\bigg{)}\\ \Longrightarrow\quad&u+\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}=\frac{h_{\nu_{u}}}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}=\operatorname{\dim_{H}}\nu_{u}.\end{split}$
Substituting this identity into (5.13) and then into (5.12), we conclude that
$\mathcal{H}_{\infty}^{u-K\varepsilon}(A)\gg e^{-S_{n}\log|G^{\prime}|(x)\cdot\operatorname{\dim_{H}}\nu_{u}}\asymp|I_{n}(\bm{a}_{n})|^{\operatorname{\dim_{H}}\nu_{u}}.$
Finally, observe that by construction,
$A\subset\left\{x\in I_{n}(\bm{a}_{n}):a_{n+1}(x)\cdots a_{n+m}(x)\geq B^{n}\right\},$
so we have the lower bound
$\mathcal{H}_{\infty}^{u-K\varepsilon}\left(\left\{x\in I_{n}(\bm{a}_{n}):a_{n+1}(x)\cdots a_{n+m}(x)\geq B^{n}\right\}\right)\geq|I_{n}(\bm{a}_{n})|^{\operatorname{\dim_{H}}\nu_{u}}.$
Applying Theorem 1.4, it follows that
$F_{m}(B)\in\mathscr{G}^{u-K\varepsilon}([0,1]).$
Since $\varepsilon>0$ is arbitrary, we conclude that
$F_{m}(B)\in\mathscr{G}^{u}([0,1]),$
which completes the proof. ∎